Consider the following time series.

t 1 2 3 4 5
yt 5 11 8 14 15
(b) Use simple linear regression analysis to find the parameters for the line that minimizes MSE for this time series.
If required, round your answers to two decimal places.
y-intercept, b0 =
Slope, b1 =
MSE =
(c) What is the forecast for t = 6?
If required, round your answer to one decimal place.
Answer #1

Let the Regression line be y = bo + b1t

where,

bo = ( Σy Σt2 - Σt Σty ) / ( nΣt2 - (Σt)2 )

b1 = ( nΣty - ΣtΣy ) / ( nΣt2 - (Σt)2 )

t y t2 ty
1 5 1 5
2 11 4 22
3 8 9 24
4 14 16 56
5 15 25 75
Total 15 53 55 182

bo = ( 53*55 - 15*182 ) / ( 5*55 - 152 ) = 3.7

b1 = ( 5*182 - 15*53 ) / ( 5*55 - 152 ) = 2.3

Hence, y = 3.7 + 2.3t

Forecasting the value using the above trend line.

MSE is the average of squared error

t y forecast error squared error
1 5 6 -1 1
2 11 8.3 2.7 7.29
3 8 10.6 -2.6 6.76
4 14 12.9 1.1 1.21
5 15 15.2 -0.2 0.04
MSE 3.26

Hence, MSE = 3.26

(c) when t = 6, forecast y = 3.7 + 2.3*6 = 17.5